# Concatenate, Fine-tuning, Re-training: A SAM-enabled Framework for Semi-supervised 3D Medical Image Segmentation

### v1

@@ Abstract
Segment Anything Model (SAM) fine-tuning has shown remarkable performance in medical image segmentation in a fully supervised manner, but requires precise annotations. To reduce the annotation cost and maintain satisfactory performance, in this work, we leverage the capabilities of SAM for establishing semi-supervised medical image segmentation models. Rethinking the requirements of effectiveness, efficiency, and compatibility, we propose a three-stage framework, i.e., Concatenate, Fine-tuning, and Re-training (CFR). The current fine-tuning approaches mostly involve 2D slice-wise fine-tuning that disregards the contextual information between adjacent slices. Our concatenation strategy mitigates the mismatch between natural and 3D medical images. The concatenated images are then used for fine-tuning SAM, providing robust initialization pseudo-labels. Afterwards, we train a 3D semi-supervised segmentation model while maintaining the same parameter size as the conventional segmenter such as V-Net. Our CFR framework is plug-and-play, and easily compatible with various popular semi-supervised methods. Extensive experiments validate that our CFR achieves significant improvements in both moderate annotation and scarce annotation across four datasets. In particular, CFR framework improves the Dice score of Mean Teacher from 29.68% to 74.40% with only one labeled data of LA dataset. The code is available at https://github.com/ShumengLI/CFR.
@@
3D medical image, segmentation, SAM-enabled, semi-supervised learning, concatenate, fine-tuning, re-training.I Introduction
Recently, general foundation models for visual segmentation [1, 2, 3, 4] have attracted widespread attention in the field of medical images owing to their excellent segmentation and generalization capabilities. Although these foundational models have made remarkable progress in medical image analysis, it is sometimes challenging to utilize a unified model to segment all medical images due to the inevitable factors, e.g., specific modalities, complex imaging techniques, and variable tissues. To tackle this issue, several recent works have been proposed to either fine-tune or add additional adapters to borrow the ability of foundation model, e.g., MSA [5] and SAMed [6] derived from Segment Anything Model (SAM) [3], to their specific tasks.
We notice that all these works are fully supervised methods with fine-tuning or adaptation techniques employed. However, fully supervised medical image segmentation relies on a great amount of precise annotations delineated by experienced experts, which makes the labeling process tedious, time-consuming, and even subjective.
@@
How to initialize effectively? According to previous studies [11, 12, 13, 14, 15, 10, 16, 9, 8, 17, 18], in the semi-supervised scenario, the quality of pseudo-labels in initialization stages plays an important role in the following segmentation. Unlike natural images, inter-slice continuity of 3D medical images is crucial for accurate target segmentation. Also, medical images usually have relatively low resolution. The existing strategy of fitting medical images, either directly enlarging 2D slices [19] or resizing the positional embeddings [6, 20], disregards the inherent inter-slice correlation that exists in 3D images. Therefore, is there a better way to boost the quality of initial pseudo-labels for medical images using a foundation model?
How to improve efficiency? Foundation model is pre-trained on a large-scale dataset whose parameter size is relatively large. Previous fine-tuning methods [19, 5] still preserve the original parameter size as foundation model during inference. During segmenting medical images, do we really need such a large parameter size? Firstly, the existing model only involving a small-size parameter indeed performs well in segmenting organs [10, 17]. Secondly, we notice that very recent works [21, 22, 23] depict the redundancy issue in the foundation model, revealing that the large-scale pre-trained models are over-parameterized. Thirdly, the appearance of medical images often has standardized views and relatively limited texture variants compared with the natural image [24]. Regarding these aforementioned issues, is it possible to escape from over-parameterized foundation model while maintaining promising results?
@@
Being aware of these observations, we believe, that in the era of foundation model, a promising semi-supervised medical image segmentation framework should be performance effectiveness, parameter efficiency, and excellent compatableness. Thus, in this paper, we propose a straightforward framework of Concatenate, Fine-tuning, and Re-training (CFR) to accomplish our above goals. We first develop a concatenate strategy, performing a concatenation operation on slices to produce an image that matches the high-resolution input, which better exploits inter-slice relationships and dimensional information. The concatenated images are then fed into the SAM’s image encoder for fine-tuning. Afterwards, we train a small-scale 3D segmentation model with the guidance of SAM while maintaining the same parameter size. The fine-tuned SAM provides a favorable initialization and is able to be compatible with various 3D segmentation models. We conduct the semi-supervised scenarios of moderate annotation111Moderate refers to a commonly used level of annotation [9, 8]. and scarce annotation222Scarce means very few annotations such as 1 labeled data. on four datasets. The experiments demonstrate our CFR framework achieves extremely close performance to full supervision with moderate annotation and exhibits remarkable improvement with scarce annotation.
@@
We propose a novel framework leveraging the ability of the foundation model while ensuring performance under semi-supervised segmentation and further reducing labeling costs, which involves three stages, i.e., concatenate, fine-tuning, and re-training.
Our concatenate strategy is simple yet effective in pseudo-labeling initialization, and largely distinct from current resize/directly fine-tuning strategy.
Our parameter size maintains the same level as the mainstream segmenter, e.g., V-Net [26], which is greatly smaller than that of foundation model.
@@
The remaining of this paper is organized as follows: Section II describes the related works on foundation models in medical images and recent semi-supervised techniques. Section III presents our proposed framework, and Sections IV, V, and VI elaborate three modules in detail. Section VII provides a summary of the framework. Section VIII reports and analyzes the experimental results, and Section IX draws the conclusion.
@@ II-A Foundation Models in Medical Images
II-A1 Vision Foundation Models
Nowadays, visual foundation models have gained significant attention and have shown impressive performance in various computer vision tasks including segmentation. Prominent examples of these models include SAM [3], SegGPT [1], SEEM [2], and SLiMe [4], along with their extended applications [27, 28, 29]. These models leverage large-scale image datasets to learn universal visual representations and demonstrate remarkable generalization ability.
In the field of medical images, UniverSeg [30] achieves universal segmentation for 2D medical images by providing an example set of image-label pairs. STU-Net [31] is a foundation model specializing in CT modalities, with its largest variant consisting of 1.4 billion parameters. Furthermore, SAM [3] has emerged as one of the most prevailing models for image segmentation, and many works extend it to medical images. SAM-Med2D [32] is a 2D model that fine-tunes SAM on 4.6 million medical images. SAM-Med3D [33] adopts a SAM-like architecture, but it is trained from scratch without utilizing the pre-trained weights of SAM.
Due to SAM’s impressive performance and broad applicability, it serves as the default foundation model in our proposed framework. SAM consists of three components: image encoder, prompt encoder, and mask decoder. The image encoder and prompt encoder extract image features and embed prompts respectively, and the mask decoder receives both embeddings and predicts the masks.
@@ II-A2 Adapt SAM to 3D Medical Images
Due to the lack of medical image data and corresponding semantic labels, SAM’s zero-shot segmentation capability is insufficient to ensure direct application in medical images [34, 35]. To extend the powerful segmentation ability to medical images, many works are devoted to fine-tuning SAM with different image processing and fine-tuning strategies.
For image processing, the disparity in image resolution between 3D medical images and pre-trained natural images poses a challenge, and two strategies have been proposed to tackle this issue. The first is upsampling fine-tuning [19], which involves upsampling each slice to match the input resolution directly. The second is small-size fine-tuning [6, 20], which reduces the input size through bilinear interpolation. However, both of these strategies are based on 2D inputs, and for 3D medical images, the predictions need to be generated by segmenting each slice. [36] and [33] extended SAM to 3D architecture, but it increases additional training overhead with a large model size. Our concatenate strategy takes into account both image dimension and image resolution differences. By creating a large-sized concatenated image, spatial information of adjacent slices could also be captured.
The fine-tuning approaches include fine-tuning only sub-parts parameters and incorporating adapters. [19] froze the image encoder and prompt encoder and only fine-tuned the mask decoder of SAM, but the performance still lags behind medical-specific models, particularly in terms of boundary consensus. [5] and [6] adopted the parameter-efficient fine-tuning techniques, with adapter strategy and low-rank-based fine-tuning strategy (LoRA) [37] respectively for fine-tuning.
@@ II-B Semi-supervised Medical Image Segmentation
Since the pixel-wise annotations require tremendous delineation time, semi-supervised learning (SSL) for segmentation aims to reduce the annotation burden by leveraging a large number of unlabeled samples along with a limited number of labeled samples [12, 13, 14, 15, 10, 16, 9, 8, 17, 18]. Semi-supervised segmentation methods mainly include pseudo-labeling and consistency regularization. Self-training [38, 39, 40] gradually generates pseudo-labels for unlabeled data through an iterative process to be jointly trained with labeled data. Mean Teacher [13] is a classic consistency regularization-based method that effectively reduces over-adaptation. Xu et al . [9] improved the classical MT model to an ambiguity-consensus mean teacher model, and Chen et al . [8] proposed a data augmentation strategy based on partition-and-recovery cubes. Existing SSL methods have demonstrated promising results with moderate annotation levels. On one hand, our framework ensures seamless integration with various SSL models. On the other hand, we also apply the framework to challenging scenarios, further reducing the annotation requirements.
@@
III Framework: CFR
III-A Notations
Formally, we now provide our notation used in this paper. Given labeled images and unlabeled images, the -th () labeled image and its ground truth are denoted as and . The -th () unlabeled image is denoted as . Here, , , , where , , indicate the corresponding dimensionality of 3D medical images and the input patch size for training is square, i.e., . is the number of different classes to segment. Our CFR framework consists of three modules as follows.
III-B Modules
Step 1: Concatenate Module. As aforementioned, by reducing the resolution difference between pre-trained samples (i.e., natural image) and fine-tuned samples (i.e., medical image), the concatenate module could transform a 3D labeled volume to a large-sized 2D image with a slice concatenation function . Also, the concatenated ground truth is obtained similarly. and are arranged in an grid with , and we concatenate zeros after all slices if .
Step 2: Fine-tuning Module. In this module, we first utilize the concatenated labeled images along with their ground truth to fine-tune a SAM parameterized by via popularly used strategy, e.g., LoRA [37, 6]. This aims to narrow the possible distribution shift between natural and medical images. Taking LoRA as an example, we denote the fine-tune function as , which is parameterized by with input as and . The updated SAM with optimal parameter is obtained as follows:
Then, we produce high-quality pseudo-labels for unlabeled images by using the prediction function of fine-tuned SAM, and generate 3D pseudo-labels by a concatenated inverse transform as follows:
@@ III-B Modules
IV Concatenate Module
To adapt from 2D natural images to 3D medical images, we recognize that the input resolution and image dimension are crucial factors. The inter-slice spatial information of 3D volumes is relevant for target recognition, and it is difficult for the large model trained at high-resolution images to generalize to low-resolution medical image slices.
@@ IV Concatenate Module
Inspired by this, our concatenate strategy, illustrated in Figure 3, matches medical images to natural image resolution, and supplements the spatial arrangement specific to 3D medical images. The input spatial resolution of the pre-trained SAM model is . Our concatenate strategy flattens the 3D volume (3D raw image or 3D patch) slice by slice and arranges them in an grid, forming a 2D image with size . Considering that the slice sizes of original medical 3D datasets might vary a lot in different cases, how to concatenate should be analyzed case by case. For the small-scale slices, such as the LA and BraTS datasets, we use a raw 3D image as an input volume. For large-scale slices, such as the BTCV and MACT datasets, we follow the common 3D processing way where dividing the volume into patches as an input volume, and then concatenate all slices to achieve the final size of , avoiding downsizing a large slice directly. Compared to the small-size input fine-tuning method [19] and the direct upsampling fine-tuning methods [6, 20] in Figure 5a, we find our concatenate strategy effectively addresses the challenges of image dimension and resolution differences. We thoroughly investigate the concatenate strategy and conduct some empirical analysis.
IV-A Slice Continuity
@@
Due to the inherent spatial continuity of 3D medical images, we explore the relationship between slice concatenation and slice order. For a concatenated 2D image, the segmentation model learns the feature correlation across slices through the self-attention mechanism, so it struggles to capture contextual information and coherence of shape without the slice order. To investigate the importance of slice continuity, we disrupt the continuity in two ways: 1) Randomly shuffling the order of the slices; 2) Randomly rotate and flip each individual slice. As illustrated in Figure 4, it has been observed that disrupting the slice order leads to the loss of shape coherence and a decrease in performance, and the results emphasize the importance of slice continuity in 3D organ segmentation.
IV-B Contextual Integrity
Our concatenate module reorganizes a 3D volume (3D raw image or 3D patch) into a 2D image, enabling a complete representation of the volume within a single image. To explore the impact of contextual integrity on concatenate fine-tuning, we approach it from two perspectives.
Firstly, we adjust the concatenation scale as shown in Figure 5b. Take the LA dataset as an example, when the concatenation scale is , the large-sized 2D image is composed of a complete 3D volume. The results indicate that there is an upward trend in overall performance as the concatenation scale increases. Specifically, the segmentation performance is best when the concatenated image contains the complete 3D volume.
@@ IV-B Contextual Integrity
Secondly, while keeping the resolution of each slice constant, we explore the impact of changing the number of slices and concatenating with natural images. As shown in Figure 4, we transition from natural images to medical images by progressively increasing the number of medical slices. The natural images are from the PASCAL VOC 2012 dataset [41]. Although this approach seems to gradually adjust from natural image features to medical image features, it actually disrupts the integrity of the anatomical structure in medical images and leads to a decrease in performance.
@@
These observation results reveal the importance of maintaining slice continuity and contextual integrity for our concatenate strategy, effectively bridging the domain and spatial dimensional gaps between natural and medical images.
V Fine-tuning Module

### v2

# Stitching, Fine-tuning, Re-training: A SAM-enabled Framework for Semi-supervised 3D Medical Image Segmentation

@@ Abstract
Segment Anything Model (SAM) fine-tuning has shown remarkable performance in medical image segmentation in a fully supervised manner, but requires precise annotations. To reduce the annotation cost and maintain satisfactory performance, in this work, we leverage the capabilities of SAM for establishing semi-supervised medical image segmentation models. Rethinking the requirements of effectiveness, efficiency, and compatibility, we propose a three-stage framework, i.e., Stitching, Fine-tuning, and Re-training (SFR). The current fine-tuning approaches mostly involve 2D slice-wise fine-tuning that disregards the contextual information between adjacent slices. Our stitching strategy mitigates the mismatch between natural and 3D medical images. The stitched images are then used for fine-tuning SAM, providing robust initialization of pseudo-labels. Afterwards, we train a 3D semi-supervised segmentation model while maintaining the same parameter size as the conventional segmenter such as V-Net. Our SFR framework is plug-and-play, and easily compatible with various popular semi-supervised methods. We also develop an extended framework SFR+ with selective fine-tuning and re-training through confidence estimation. Extensive experiments validate that our SFR and SFR+ achieve significant improvements in both moderate annotation and scarce annotation across five datasets. In particular, SFR framework improves the Dice score of Mean Teacher from 29.68% to 74.40% with only one labeled data of LA dataset. The code is available at https://github.com/ShumengLI/SFR.
@@
3D medical image segmentation, semi-supervised learning, stitching, fine-tuning, re-training, SAM-enabled.I Introduction
Recently, general foundation models for visual segmentation [1, 2, 3, 4] have attracted widespread attention in the field of medical images owing to their excellent segmentation and generalization capabilities. Although these foundational models have made remarkable progress in medical image analysis, it is sometimes challenging to utilize a unified model to segment all medical images due to the inevitable factors, e.g., specific modalities, complex imaging techniques, and variable tissues. To tackle this issue, several recent works have been proposed to either focus on prompt engineering [5, 6, 7] or design adapters for fine-tuning [8, 9, 10, 11, 12, 13, 14] to borrow the ability of foundation model, e.g., MSA [9] and SAMed [10] derived from SAM [3], to their specific tasks.
We notice that most of these works [8, 9, 10, 11, 13, 14, 5] are fully supervised methods with fine-tuning or adaptation techniques employed. However, fully supervised medical image segmentation relies on a great amount of precise annotations delineated by experienced experts, which makes the labeling process tedious, time-consuming, and even subjective.
@@
How to initialize effectively? According to previous studies [19, 20, 21, 22, 23, 24, 18, 17, 16, 25], in the semi-supervised scenario, the quality of pseudo-labels in initialization stages plays an important role in the following segmentation. Unlike natural images, inter-slice continuity of 3D medical images is crucial for accurate target segmentation. Also, medical images usually have relatively low resolution. The existing strategies [8, 10, 11, 12, 26] of fitting medical images, including directly enlarging 2D slices [8] and resizing the positional embeddings [10, 11], use slice by slice fine-tuning and disregard the inherent inter-slice correlation that exists in 3D images. Therefore, is there a better way to boost the quality of initial pseudo-labels for medical images using a foundation model?
How to improve efficiency? Foundation model is pre-trained on a large-scale dataset whose parameter size is relatively large. Existing fine-tuning methods [8, 9, 12, 13, 14, 5, 6, 7] still preserve the original parameter size as foundation model during inference or even introduce additional parameters. During segmenting medical images, do we really need such a large parameter size? Firstly, the existing model only involving a small-size parameter indeed performs well in segmenting organs [18, 25]. Secondly, we notice that very recent works [27] depict the redundancy issue in the foundation model, revealing the large-scale pre-trained models are over-parameterized. Thirdly, the appearance of medical images often has standardized views and relatively limited texture variants compared with natural images [28]. Regarding these issues, is it possible to escape from over-parameterized foundation model while maintaining promising results?
@@
Being aware of these observations, we believe, that in the era of foundation model, a promising semi-supervised medical image segmentation framework should be performance effectiveness, parameter efficiency, and excellent compatableness. Thus, we propose a straightforward framework of Stitching, Fine-tuning, and Re-training (SFR) to accomplish our above goals. We first develop a stitching strategy, performing a stitching operation on slices to produce an image that matches the high-resolution input, which better exploits inter-slice relationships and dimensional information. The stitched images are then fed into the SAM for fine-tuning. Afterwards, we train a small-scale 3D segmentation model with the guidance of SAM while maintaining the same parameter size. The fine-tuned SAM provides a favorable initialization and is compatible with various 3D models. In addition, we develop the extended framework SFR+, introducing confidence estimation and selective training strategy to enhance the utilization of unlabeled data. We conduct the semi-supervised scenarios of moderate annotation111Moderate refers to a commonly used level of annotation [17, 16]. and scarce annotation222Scarce means very few annotations such as 1 labeled data. on five datasets. The experiments demonstrate our SFR and SFR+ frameworks achieve extremely close performance to full supervision with moderate annotation and exhibit remarkable improvement with scarce annotation. As shown in Fig. 1, our method in fine-tuning stage outperforms other extended foundation models and achieves a large improvement in re-training stage.
@@
We propose a novel framework leveraging the ability of the foundation model while ensuring performance under semi-supervised segmentation and further reducing labeling costs, which involves three stages, i.e., stitching, fine-tuning, and re-training.
Our stitching strategy is simple yet effective in pseudo-labeling initialization, and largely distinct from current resize/directly fine-tuning strategy.
Our parameter size during inference maintains the same level as the mainstream segmenter, e.g., V-Net [30], which is greatly smaller than that of foundation model.
@@ II-A Foundation Models in Medical Images
II-A1 Visual Foundation Models
Nowadays, visual foundation models have gained significant attention and have shown impressive performance in various computer vision tasks including segmentation. Prominent examples of these models include SAM [3], SegGPT [1], SEEM [2], SLiMe [4], and SAM 2 [31], along with their extended applications [32, 33, 34]. These models leverage large-scale image datasets to learn universal visual representations and demonstrate remarkable generalization ability.
In the field of medical images, UniverSeg [35] achieves universal segmentation for 2D medical images by providing an example set of image-label pairs. STU-Net [36] is a foundation model specializing in CT modalities, with its largest variant consisting of 1.4 billion parameters. Furthermore, SAM [3] has emerged as one of the most prevailing models for image segmentation, and many works extend it to medical images. SAM-Med2D [37] is a 2D model that fine-tunes SAM on 4.6 million medical images. SAM-Med3D [38] adopts a SAM-like architecture, but it is trained from scratch without utilizing the pre-trained weights of SAM. Due to SAM’s impressive performance and broad applicability, it serves as the default foundation model in our proposed framework.
@@ II-A2 Adapt SAM to 3D Medical Images
SAM’s zero-shot capability is insufficient to ensure direct application in medical images [39, 40]. To extend the powerful segmentation ability to medical images, many works are devoted to fine-tuning with different image processing and fine-tuning strategies.
For image processing, the disparity in image resolution between 3D medical images and pre-trained natural images poses a challenge, and two strategies have been proposed to tackle this issue. The first is upsampling fine-tuning [8], which involves upsampling each slice to match the input resolution directly. The second is small-size fine-tuning [10, 11], which reduces the input size through bilinear interpolation. However, both of these strategies are based on 2D inputs, and for 3D medical images, the predictions need to be generated by segmenting each slice. 3DSAM-Adapter [41] and SAM-Med3D [38] extend SAM to 3D architecture, whereas they increase additional training overhead with a large model size. In contrast, our stitching strategy is designed to accommodate variations in image dimension and resolution, creating large-sized stitched images that capture spatial information across adjacent slices effectively.
The fine-tuning approaches include fine-tuning only sub-parts parameters and incorporating adapters. MedSAM [8] fine-tunes the mask decoder of SAM and freezes the encoders, whereas the performance shows a lag behind medical-specific models, particularly in terms of boundary area. MSA [9] and SAMed [10] adopt the parameter-efficient fine-tuning techniques, using adapter and low-rank-based strategy (LoRA) [42] strategies for fine-tuning.
@@ II-B Semi-supervised Medical Image Segmentation
Since the pixel-wise annotations require tremendous delineation time, semi-supervised learning (SSL) for segmentation aims to reduce the annotation burden by leveraging a large number of unlabeled samples along with a limited number of labeled samples [20, 21, 22, 23, 24, 18, 17, 16, 25]. Semi-supervised segmentation methods mainly include pseudo-labeling and consistency regularization. Self-training [43] gradually generates pseudo-labels for unlabeled data through an iterative process to be jointly trained with labeled data. Mean Teacher [21] is a classic consistency regularization-based method that effectively reduces over-adaptation. Xu et al . [17] improves the classical MT model to an ambiguity-consensus mean teacher model, and Chen et al . [16] develops a data augmentation strategy based on partition-and-recovery cubes. Existing SSL methods have demonstrated promising results with moderate annotation levels. On one hand, our framework ensures seamless integration with various SSL models. On the other hand, we also apply the framework to challenging scenarios, further reducing the annotation requirements.
Some of the latest works [26, 7] explore incorporating SAM into semi-supervised training. For example, ASLseg [26] couples SAM with a specific semi-supervised model to refine pseudo-labels, while CPC-SAM [7] uses SAM prompts for cross-branch teaching. However, these methods still rely on slice-based processing and preserve a high computational load during inference. Our method could help improve performance while reducing inference costs. It is worth mentioning that our SFR and SFR+ frameworks are compatible with various semi-supervised segmentation methods for medical images.
@@
III Method
III-A Notations and Framework Overview
Formally, we now provide our notation used in this paper. Given labeled images and unlabeled images, the -th () labeled image and its ground truth are denoted as and . The -th () unlabeled image is denoted as . Here, , , , where , , indicate the corresponding dimensionality of 3D medical images and the input patch size for training is square, i.e., . is the number of different classes to segment.
As illustrated in Fig. 2, we build our SFR framework, which consists of the following three modules: the Stitching Module, the Fine-tuning Module, and the Re-training Module. The stitching module mitigates the mismatch between natural and 3D medical images. The stitched images are input into SAM to fine-tune and provide initial pseudo labels for the semi-supervised module. Afterwards, the 3D semi-supervised segmentation model is trained. Furthermore, we enhance our framework by proposing SFR+, which selectively fine-tunes and re-trains through confidence estimation.
Step 1: Stitching Module. As aforementioned, by reducing the resolution difference between pre-trained samples (i.e., natural image) and fine-tuned samples (i.e., medical image), the stitching module could transform a 3D labeled volume to a large-sized 2D image with a slice stitching function . Also, the stitched ground truth is obtained similarly. and are arranged in a grid with , and we stitch zeros after all slices if .
Step 2: Fine-tuning Module. We first utilize the stitched labeled images along with their ground truth to fine-tune a SAM parameterized by via popularly used strategy, e.g., LoRA [42, 10]. This aims to narrow the possible distribution shift between natural and medical images. Taking LoRA as an example, we denote the fine-tune function as , which is parameterized by with input as and . The updated SAM with optimal parameter is obtained as follows:
Then, we produce high-quality pseudo-labels for unlabeled images by using the prediction function of fine-tuned SAM, and generate 3D pseudo-labels by a stitching inverse transform as follows:
@@ III-A Notations and Framework Overview
III-B SFR Framework
III-B1 Stitching Module
@@
To adapt from 2D natural images to 3D medical images, we recognize that the input resolution and image dimension are crucial factors. The inter-slice spatial information of 3D volumes is relevant for target recognition, and it is difficult for the large model trained at high-resolution images to generalize to low-resolution medical image slices. Inspired by this observation, our stitching strategy, illustrated in Fig. 3, matches medical images to natural image resolution, and supplements the spatial arrangement specific to 3D medical images. The input spatial resolution of the pre-trained SAM model is . Our stitching strategy arranges the 3D volume (either a raw 3D image or a 3D patch) slice by slice into a grid, producing a 2D image of size . Regarding the variability in slice sizes across different medical datasets, there is a performance trade-off between the slice resolution and the number of stitched slices. Our method could effectively manage images with different numbers of slices. For the small-scale slices, such as the LA [15] dataset, we use a raw 3D image as an input volume. For large-scale slices, such as the BTCV [44] dataset, we follow the common 3D processing way where dividing the volume into patches as an input volume, and then stitch all slices to achieve the final size of , avoiding downsizing a large slice directly. Compared to the small-size input fine-tuning method [10, 11] and the direct upsampling fine-tuning methods [8] in Fig. 3, we find our stitching strategy effectively addresses the challenges of image dimension and resolution differences.
We thoroughly investigate the stitching strategy from slice continuity and contextual integrity.
@@
Slice Continuity. Due to the inherent spatial continuity of 3D medical images, we explore the relationship between slice stitching and slice order. For a stitched 2D image, the segmentation model learns the feature correlation across slices through the self-attention mechanism, so it struggles to capture contextual information and coherence of shape without the slice order. To investigate the importance of slice continuity, we disrupt the continuity in two ways: 1) Randomly shuffling the order of the slices; 2) Randomly rotate and flip each individual slice. The same slicing operation is also applied to the ground truth masks to maintain consistency between the input data and labels. As illustrated in Fig. 4, it has been observed that disrupting the slice order leads to the loss of shape coherence and a decrease in performance, and the results emphasize the importance of slice continuity in 3D organ segmentation.
@@
Contextual Integrity. Our stitching module reorganizes a volume (3D raw image or 3D patch) into a image, enabling a complete representation of the volume within a single image. To explore the impact of contextual integrity on stitching, we stitch the medical slices with natural images while keeping the resolution of each slice constant. As shown in Fig. 5, we transition from natural images to medical images by progressively increasing the number of medical slices. Specifically, at the beginning of the training process, we incorporate natural images from the PASCAL VOC 2012 dataset [45], which contains a total of 2,913 images. For each iteration, we randomly select a batch of natural images from this dataset. As training progresses, we incrementally replace parts of the natural images with medical image slices, such that by the final stage, only medical slices are utilized. For testing, we consistently use fully stitched medical image slices. Although this approach seems to gradually adjust from natural image features to medical image features, it actually disrupts the integrity of the anatomical structure in medical images and leads to a decrease in performance.
@@
These observation results reveal the importance of maintaining slice continuity and contextual integrity for our stitching strategy, effectively bridging the domain and spatial dimensional gaps between natural and medical images.
III-B2 Fine-tuning Module